OpenStudy (anonymous):

Which statement is false? A. The number zero is rational number. B. Some irrational numbers are also rational numbers. C. Every irrational number is a real number. D. Every integer is a rational number

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

Im thinking of c. idk maybe d

OpenStudy (anonymous):

what are yours?

jimthompson5910 (jim_thompson5910):

What is an integer?

OpenStudy (anonymous):

A number

jimthompson5910 (jim_thompson5910):

what kind of number?

jimthompson5910 (jim_thompson5910):

is something like 2.5 an integer?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

why not?

OpenStudy (anonymous):

2.5 is a Rational Number

OpenStudy (anonymous):

No wait

jimthompson5910 (jim_thompson5910):

yes it is, do you see why?

OpenStudy (anonymous):

because it can be written as a fraction

jimthompson5910 (jim_thompson5910):

a fraction of what type of number?

OpenStudy (anonymous):

Rational Number

jimthompson5910 (jim_thompson5910):

a rational number is a fraction of _______ (fill in the blank)

OpenStudy (anonymous):

Natural Number?

jimthompson5910 (jim_thompson5910):

close, but you can use negative numbers too

jimthompson5910 (jim_thompson5910):

so it's actually integers

jimthompson5910 (jim_thompson5910):

for instance, 2/3 is a rational number since 2 and 3 are integers

OpenStudy (anonymous):

Oh ok sorry im kinda new to this

jimthompson5910 (jim_thompson5910):

thats fine

OpenStudy (anonymous):

im kinda dumb anyway so

jimthompson5910 (jim_thompson5910):

well you just need more practice. You aren't dumb

jimthompson5910 (jim_thompson5910):

saying "Every integer is a rational number" is true because we can turn an integer like 7 into 7/1 or an integer like 22 into 22/1 etc etc

jimthompson5910 (jim_thompson5910):

they want to know which statement is false, so we can cross off choice D

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i have another question.

jimthompson5910 (jim_thompson5910):

whats that

OpenStudy (anonymous):

Which of these numbers can be classified as both real and irrational? A. 5.85 B. 63.4 C. Square root of 21 D. Square root of 36

jimthompson5910 (jim_thompson5910):

well you didn't finish with the last question

jimthompson5910 (jim_thompson5910):

but anything irrational is something that CANNOT be written as a fraction of integers

jimthompson5910 (jim_thompson5910):

so it's the complete opposite of something like 2/3

OpenStudy (anonymous):

oh wait i think i have it. isnt it B. 63.4

jimthompson5910 (jim_thompson5910):

Rule: all terminal decimals can be written as a fraction ex: 2.5 = 25/10 = 5/2

jimthompson5910 (jim_thompson5910):

fraction of two integers I meant

jimthompson5910 (jim_thompson5910):

so all terminal decimal numbers are rational

jimthompson5910 (jim_thompson5910):

Another Rule: repeating decimal numbers are also rational There has to be a clear pattern repeating forever

jimthompson5910 (jim_thompson5910):

example: 2/11 = 0.181818181818181818... notice how the "18"s repeat forever